# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ E.1 Runtime complexity of Algorithm˜1 by using LMC
$L\times n\geq\tilde{\mathcal{O}}\left(\frac{d\texttt{k}\mathrm{B}^{2}\log(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right).$
@@
$L\;=\;\mathcal{O}\bigl(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log(\Delta^{-1})\bigr).$
@@
$L=\mathcal{O}\bigl(\texttt{k}\,\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log(\epsilon_{v}^{-1})\bigr).$ (83)
@@ 3. This is exactly (84):
$L\times n\;=\;\tilde{\mathcal{O}}\!\left(\frac{d\,\texttt{k}\,\mathrm{B}^{2}\log(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right),$
@@
First, we change Algorithm˜2 to algorithm˜4 by replacing the samples of Gibbs measure with samples of the output of parallel LMCs and changing PSGD to PSGD-LMC (see Algorithm˜3) in lines 4 and 5. To apply LMC for simulating the Gibbs measure, based on the proof of ˜5.5, the only part we must change is to bound the last term in (73). What we need is to change the bound in (74) on the last of term of right-hand side in (73) for samples from Gibbs to samples from the outputs of parallel LMCs. We provide the proof of ˜5.7 in the following:
@@
Step 1: Zeroth-order analysis (same as Theorem 5.5). As in the proof of ˜5.5, we work with the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\mathrm{Unif}(\mathbb{S}^{m-1})}\big[F_{\max}(\theta+\rho u)\big].$ Using the smoothness of $\psi_{\rho}$ and the projected gradient-mapping update, the standard analysis of constrained zeroth-order methods yields (see eq.˜73 in Section˜D.3)
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}\big]\;\leq\;\frac{8\big(\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]\big)}{\tilde{\eta}_{i}}\;+\;28\,\mathbb{E}\big[\|\mathsf{g}_{\rho}(\theta_{i})-\hat{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big],$ (92)
@@
and $\hat{\mathsf{g}}_{\rho}(\theta_{i})$ is the stochastic gradient estimator used in Algorithm 4. Up to (92), the analysis is identical to Theorem 5.5; the only new task is to control $\mathbb{E}[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta]$ when $\hat{\mathsf{g}}_{\rho}$ is built from the LMC-based oracle $\hat{\psi}$.
@@
$\mathbb{E}\big[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta\big]$
$\leq\;\frac{m^{2}}{\rho^{2}}\,\mathbb{E}\Big[\big|\hat{\psi}(\theta+\rho u_{t})-F_{\max}(\theta+\rho u_{t})\big|^{2}+\big|\hat{\psi}(\theta-\rho u_{t})-F_{\max}(\theta-\rho u_{t})\big|^{2}\Bigm|\theta\Big].$ (94)
@@
Step 3: Convergence rate and choice of parameters. Plugging (98) into (92), choosing a constant stepsize $\tilde{\eta}_{i}=\eta\leq m^{-1/2}\rho L_{F_{\max},1}^{-1}$, and summing $i=0,\dots,N-1$ exactly as in the proof of Theorem 5.5 gives
$\frac{1}{N}\sum_{i=0}^{N-1}\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\eta)\|^{2}\big]\;=\;\mathcal{O}\!\left(\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{\rho^{2}}\epsilon_{v}^{2}\right).$
@@
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho}(\hat{\theta});\eta)\|\big]=\mathcal{O}\!\left(\sqrt{\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}}+\frac{m}{\rho}\epsilon_{v}\right).$
Setting $\epsilon_{v}:=\rho\epsilon m^{-1}$ and taking $N\gtrsim m\,\mathrm{diam}(\Theta)\,\rho^{-2}\epsilon^{-2}$ yields
@@
Each outer iteration of Algorithm 4 uses $b_{u}$ two-point evaluations of the oracle $\hat{\psi}$. Each call to $\hat{\psi}$ in turn runs PSGD-LMC with $|I|$ parallel LMC chains of length $n$. Thus the total number of lower-level gradient evaluations is
$N\times n\times\bigl(b_{u}L+b_{u}|I|\bigr).$
Substituting the choices of $N$, $n$, $L$, and $|I|$ from Lemma E.1 and the LMC mixing bound (Proposition 5.6), with the choice $\epsilon_{v}=\rho\epsilon m^{-1}$, gives
@@
$N\times n\times(b_{u}L+b_{u}|I|)=\mathcal{O}\!\left(\frac{d\,m^{9+8\texttt{k}}\,[\texttt{k}\,\mathrm{B}^{2}+\tilde{C}_{f}]\log(\epsilon^{-1})}{\epsilon^{8\texttt{k}+10}\rho^{8\texttt{k}+10}}\right),$
which is the claimed complexity bound. ∎

### v4

@@ E.1 Runtime Complexity of Algorithm˜1 by Using LMC
$L\times n\geq\tilde{\mathcal{O}}\left(\frac{d\texttt{k}^{2}\mathrm{B}^{2}\log^{2}(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right).$
@@
$L\;=\;\mathcal{O}\bigl(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log^{2}(\Delta^{-1})\bigr).$
@@
$L=\mathcal{O}\bigl(\texttt{k}^{2}\,\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log^{2}(\epsilon_{v}^{-1})\bigr).$ (85)
@@ 3. This is exactly (86):
$L\times n\;=\;\tilde{\mathcal{O}}\!\left(\frac{d\,\texttt{k}^{2}\,\mathrm{B}^{2}\log^{2}(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right),$
@@
First, we change Algorithm˜2 to Algorithm˜4 by replacing the samples of Gibbs measure with samples of the output of parallel LMCs and changing PSGD to PSGD-LMC (see Algorithm˜3) in lines 4 and 5. Relative to the proof of ˜5.5, the directional-variance term in (D.3) is unchanged; the only new ingredient is the bound on the oracle term $\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\hat{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$. We provide the proof of ˜5.7 in the following:
@@
Step 1: Zeroth-order analysis (same as Theorem 5.5). As in the proof of ˜5.5, we work with the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\mathrm{Unif}(\mathbb{S}^{m-1})}\big[F_{\max}(\theta+\rho u)\big].$ Using the smoothness of $\psi_{\rho}$ and the projected gradient-mapping update, the standard analysis of constrained zeroth-order methods yields (see section˜D.3 in Section˜D.3)
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}\big]$ $\leq\frac{8\big(\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]\big)}{\tilde{\eta}_{i}}$
$\qquad\qquad+24\,\mathbb{E}\big[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big]$
$\qquad\qquad+40\,\mathbb{E}\big[\|\mathsf{g}_{\rho}(\theta_{i})-\hat{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big],$ (94)
@@
and $\hat{\mathsf{g}}_{\rho}(\theta_{i})$ is the stochastic gradient estimator used in Algorithm 4. Up to (E.2), the analysis is identical to Theorem 5.5; in particular,
$\mathbb{E}\big[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big]\leq\frac{m^{2}L_{F_{\max},1}^{2}}{b_{u}}.$
The only new task is to control $\mathbb{E}[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta]$ when $\hat{\mathsf{g}}_{\rho}$ is built from the LMC-based oracle $\hat{\psi}$.
@@
$\mathbb{E}\big[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta\big]\leq$
$\;\frac{m^{2}}{\rho^{2}}\,\mathbb{E}\Big[\big|\hat{\psi}(\theta+\rho u_{t})-F_{\max}(\theta+\rho u_{t})\big|^{2}+\big|\hat{\psi}(\theta-\rho u_{t})-F_{\max}(\theta-\rho u_{t})\big|^{2}\Bigm|\theta\Big].$ (96)
@@
Step 3: Convergence rate and choice of parameters. Plugging (100) into (E.2), choosing a constant stepsize $\tilde{\eta}_{i}=\eta\leq m^{-1/2}\rho L_{F_{\max},1}^{-1}$, and summing $i=0,\dots,N-1$ exactly as in the proof of Theorem 5.5 gives
$\frac{1}{N}\sum_{i=0}^{N-1}\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\eta)\|^{2}\big]\;=\;\mathcal{O}\!\left(\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\epsilon_{v}^{2}\right).$
@@
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho}(\hat{\theta});\eta)\|\big]=\mathcal{O}\!\left(\sqrt{\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}}+\frac{m}{\sqrt{b_{u}}}+\frac{m}{\rho}\epsilon_{v}\right).$
Setting $\epsilon_{v}:=\rho\epsilon m^{-1}$ and taking $N\gtrsim m\,\mathrm{diam}(\Theta)\,\rho^{-2}\epsilon^{-2}$ and $b_{u}\gtrsim m^{2}\epsilon^{-2}$ yields
@@
Each outer iteration of Algorithm 4 uses $b_{u}$ two-point evaluations of the oracle $\hat{\psi}$. Each call to $\hat{\psi}$ in turn runs PSGD-LMC with $|I|$ parallel LMC chains of length $n$. Consequently, the total number of lower-level gradient evaluations is $N\times n\times\bigl(b_{u}L+b_{u}|I|\bigr),$ and the total number of upper-level function evaluations is $N\times b_{u}\times|I|$. Substituting the choices of $N$, $n$, $L$, and $|I|$ from Section˜E.1 and the LMC mixing bound (Section˜5.3), with the choice $\epsilon_{v}=\rho\epsilon m^{-1}$, gives
@@
$N\times n\times(b_{u}L+b_{u}|I|)$ $=\mathcal{O}\!\left(\frac{d\,m^{11+8\texttt{k}}\,[\texttt{k}^{2}\,\mathrm{B}^{2}+\tilde{C}_{f}]\log^{2}(\epsilon^{-1})}{\epsilon^{8\texttt{k}+12}\rho^{8\texttt{k}+10}}\right),$
$N\times b_{u}\times|I|$ $=\mathcal{O}\left(\frac{m^{2\texttt{k}+5}}{\rho^{2\texttt{k}+4}\epsilon^{2\texttt{k}+6}}\right),$
which is the claimed complexity bounds. ∎
Appendix F Compute-matched Comparisons under a Fixed Wall-clock Budget
This appendix complements the iteration-based results in Section˜6 by comparing methods under a fixed wall-clock time budget. As emphasized in Section˜6, PSZO-MinSel invokes a Gibbs–superquantile oracle implemented via PSGD and parallel-chain Langevin Monte Carlo, so the cost per outer iteration can differ substantially from gradient-based baselines.
Timing protocol and selection. Toy-example methods are implemented in JAX, while data hyper-cleaning methods are implemented in PyTorch; all experiments are executed on the same machine described in Section˜6. To obtain reliable timings, we exclude one-time compilation by running a short warm-up pass before timing and synchronize the device when recording timestamps. Runtimes are measured as wall-clock seconds elapsed from the start of the timed run; for JAX experiments we time a JIT-compiled step function and record timestamps with device synchronization to avoid asynchronous timing artifacts. As in Section˜6, for each method we select the best-so-far iterate within a budget $T$, $\hat{\theta}(T)\in\arg\min\{\mathcal{E}_{n}:\ t_{n}\leq T\},$ where $\mathcal{E}_{n}$ is $F(\theta_{n})$ for the toy example and the validation loss in (34) for data hyper-cleaning. All tables below report performance evaluated at $\hat{\theta}(T)$.
F.1 Toy example: fixed-runtime results
Per-iteration timing. Table˜5 reports the mean wall-clock time per outer iteration (post warm-up) for the toy example. This highlights the substantial per-iteration cost gap between PSZO-MinSel and gradient-based baselines.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
2 0.022 0.018 0.026 0.027 6.74
5 0.021 0.018 0.027 0.032 7.90
10 0.022 0.018 0.036 0.045 8.10
20 0.021 0.017 0.045 0.046 9.70
Table 5: Mean wall-clock time per outer iteration (ms) on the toy example with $\texttt{k}=d-1$.
F.1.1 Optimistic formulation: dimension sweep under fixed time
Table˜6 replicates Table˜1 under a fixed wall-clock budget $T_{\mathrm{toy}}=10$ seconds and reports the best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
[…]
Table 6: Compute-matched analogue of Table˜1: best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$ under a fixed wall-clock budget $T_{\mathrm{toy}}=10$ s.
Under equal time, PSZO-MinSel retains a clear advantage across dimensions, consistent with the iteration-based conclusions in Section˜6.
F.1.2 Optimistic formulation: constant manifold dimension under fixed time
Table˜7 replicates Table˜2 under the same budget $T_{\mathrm{toy}}=10$ seconds, with fixed manifold dimension $\texttt{k}=1$.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
[…]
Table 7: Compute-matched analogue of Table˜2: best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$ under $T_{\mathrm{toy}}=10$ s, with fixed $\texttt{k}=1$.
With fixed intrinsic dimension, PSZO-MinSel remains stable as $d$ increases, reinforcing that performance is governed primarily by the manifold dimension rather than the ambient dimension.
F.2 Data hyper-cleaning: fixed-runtime results
Per-iteration timing.
Table˜8 reports the mean wall-clock time per outer iteration (post warm-up) on MNIST for representative models.
Model V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
Linear 0.015 0.016 15.6 0.068 13.5
2-layer MLP 0.022 0.026 18.4 0.087 17.1
Table 8: Mean wall-clock time per outer iteration (s) on MNIST data hyper-cleaning (post warm-up).
RHG and PSZO-MinSel have much higher per-outer costs in this task, making fixed-runtime evaluation particularly important.
Fixed-runtime performance.
Table˜9 replicates Table˜4 under a fixed budget $T_{\mathrm{hc}}=1800$ seconds. For each method, we select the iterate achieving the lowest validation loss within $T_{\mathrm{hc}}$ and report the corresponding test accuracy and F1-score.
Table 9: Compute-matched analogue of Table˜4 under a fixed wall-clock budget $T_{\mathrm{hc}}=1800$ seconds (pollute rate $p\in\{40,60,80\}\%$): test accuracy (%) and F1 score (%), evaluated at the best-so-far iterate within the time budget (mean $\pm$ 95% CI over 10 runs).
$p$ Method Linear model 2-layer MLP
Test acc. F1 Test acc. F1
40 RHG $60.50\pm 0.32$ $45.5\pm 0.4$ $60.54\pm 0.29$ $42.3\pm 0.4$
IAPTT-GM $\mathbf{90.4\pm 0.45}$ $80.5\pm 0.5$ $\mathbf{93.76\pm 0.33}$ $76.9\pm 0.3$
G-PBGD $83.55\pm 0.57$ $71.3\pm 0.4$ $82.64\pm 0.62$ $72.3\pm 0.4$
V-PBGD $89.43\pm 0.23$ $\mathbf{83.2\pm 0.3}$ $86.56\pm 0.10$ $\mathbf{80.2\pm 0.2}$
PSZO-MinSel $88.5\pm 0.51$ $78.2\pm 0.4$ $90.3\pm 0.44$ $79.2\pm 0.2$
60 RHG $62.74\pm 0.69$ $33.8\pm 0.4$ $63.34\pm 0.42$ $35.5\pm 0.3$
IAPTT-GM $79.42\pm 0.38$ $45.4\pm 0.3$ $78.82\pm 0.45$ $50.5\pm 0.4$
G-PBGD $69.48\pm 0.42$ $40.2\pm 0.2$ $66.65\pm 0.42$ $46.5\pm 0.6$
V-PBGD $73.14\pm 0.25$ $48.7\pm 0.3$ $72.63\pm 0.28$ $52.4\pm 0.3$
PSZO-MinSel $\mathbf{80.50\pm 0.30}$ $\mathbf{57.3\pm 0.4}$ $\mathbf{82.50\pm 0.40}$ $\mathbf{56.5\pm 0.4}$
80 RHG $50.2\pm 0.4$ $24.5\pm 0.4$ $54.5\pm 0.5$ $20.9\pm 0.3$
IAPTT-GM $65.4\pm 0.54$ $30.1\pm 0.4$ $64.54\pm 0.42$ $32.9\pm 0.2$
G-PBGD $67.5\pm 0.33$ $30.9\pm 0.8$ $62.60\pm 0.34$ $30.4\pm 0.4$
V-PBGD $64.23\pm 0.22$ $32.5\pm 0.2$ $61.22\pm 0.23$ $30.9\pm 0.4$
PSZO-MinSel $\mathbf{70.34\pm 0.45}$ $\mathbf{36.7\pm 0.5}$ $\mathbf{68.43\pm 0.38}$ $\mathbf{39.1\pm 0.4}$
At low noise ($p=40\%$), PSZO-MinSel is slightly below the strongest baselines but remains comparable under the same runtime budget; as the pollute rate increases ($p=60\%,80\%$), PSZO-MinSel achieves the best performance, consistent with the benefits of trajectory-based selection under heavier corruption.